# Long Chain-of-Thought for Reasoning Large Language Models

### v3

@@ 5.1.2 Process Feedback
Recent studies highlight the significance of feedback in developing effective PRMs for complex reasoning tasks, particularly in a step-level view [106, 342, 415]. (1) Process Annotated PRM Training: Earlier, Lightman et al. [358] demonstrate that training process feedback with human-annotated data (PRM800K) surpasses outcome supervision in creating reliable reward models. However, this approach requires significant human effort. To address this, Wang et al. [634] introduce Math-Shepherd, a dataset that generates step-by-step supervision using a Tree Search-inspired method [62, 794]. Following this, methods like QwQ [581], Skywork-o1 [450], AceMath [394], and PRIME [115] adopt similar techniques to enhance PRM performance. Additionally, Zhang et al. [826] propose entropy regularization to improve model convergence. Rather than focusing solely on the first error step, Full-Step-DPO [716] assigns rewards for the entire reasoning chain, including error steps. VersaPRM [807] extends PRMs across multiple domains, broadening their applicability. Similarly, Gu et al. [172] and Zhang et al. [852] suggest training models with student preferences aligned to teacher preferences, ensuring effective preference distillation. Further, Wang et al. [644] propose VisualPRM400K and expand this paradigm to multimodal scenarios. (2) Outcome Annotated PRM Training: Alternative approaches, such as OVM [776], Implicit PRM [793], AutoPSV [398], and DVO [827], leverage outcome supervision or implicit feedback to train PRMs, reducing the need for extensive human-annotated data [706, 513]. UAS [778] incorporates uncertainty-aware value models [215] into feedback predictions. Additionally, Aurora [561] utilizes ensemble prompting strategies and reference answers for reverse verification, training stronger PRMs that better align with the Long CoT data distribution. Furthermore, PAV [519] suggests that rewards should reflect reasoning progress, as measured by changes in the likelihood of producing a correct future response before and after each step. Yang et al. [738], Lee et al. [302], Yoon et al. [773] extend these paradigms to the token level. Further Liu et al. [389] propose AdaptiveStep for dividing the reasoning trajectory by model confidence. The resulting PRM allows for more accurate token-level evaluations compared to previous PRMs. Moreover, Chen et al. [94] expand these into interactive agent scenarios, allowing for automatically learning reward models from the environment without manual annotation.
@@
As PRM training remains heavily dependent on manually annotated data, recent research has explored methods for enabling models to generate their natural language feedback to optimize performance [722]. These approaches fall into two primary categories: (1) Model-Driven Feedback Reasoning: Earlier work such as React [756] and Reflexion [530] enhances RLLMs with natural language feedback at each action and reasoning step [152, 107, 77], improving decision-making in diverse tasks. Similarly, Step-DPO [292] uses RLLM to self-verify step-level positive and negative pairs for training through the DPO paradigm, achieving strong performance. Additionally, Sun et al. [553] propose a dynamic error classification framework that adapts based on model outputs, improving performance in mathematical reasoning tasks by addressing specific error patterns in math word problems. Furthermore, Xie et al. [704] and He et al. [193] iteratively apply MCTS to collect preference data, utilizing its forward-looking capabilities to decompose instance-level rewards into more precise step-level signals, thereby enhancing feedback accuracy. However, step-wise feedback often suffers from reliability issues, which can be mitigated by uncertainty quantification [771, 768], improving the reliability of step-wise verification in reward models for mathematical reasoning tasks. Moreover, Fu et al. [145] define the CoT Average Causal Effect (CACE) to capture causal relationships between steps, resulting in a causalized Long CoT where all steps are both correct and comprehensible. (2) Environment-Driven Feedback Reasoning: Given the increasing complexity of large models, there is growing interest in combining prompt-based LLMs with external environments to generate more interpretable and controllable feedback [700]. For example, ORPS [790] and Drori et al. [129] minimize dependence on human annotations by using execution feedback, enabling models to autonomously refine their solutions. Additionally, Shrestha et al. [531] contribute by translating model outputs into Python code, helping to identify logical errors, gain insights into flawed reasoning processes, and guide improvements in mathematical reasoning. Xu et al. [711] integrate reasoning models with an interactive environment, enabling learning in more dynamic scenarios and creating a more generalizable self-learning framework.
@@ 5.2.1 Prompt-based Refinement Generation
Research on prompt-based refine generation focuses on enhancing the performance of LLMs through iterative self-refinement mechanisms [458, 864, 82, 374, 820, 604, 654]. A prominent approach involves prompting RLLMs to generate initial outputs, followed by self-feedback that iteratively refines and improves performance across tasks such as dialogue generation and mathematical reasoning [515, 425, 870, 530, 434, 274, 600, 383], which even much reduce the hallucinations [225, 247]. Noteworthy methods, like Self-Backtracking [748], Refiner [469], and BackMath [839], allow LLMs to adjust their reasoning autonomously, reducing unnecessary complexity in decision-making [688]. Further, Havrilla et al. [189] extend the paradigm by integrating overall-level and step-level refinements, improving refinement performance. Yang et al. [751] propose a method to decompose the self-correction capability of LLMs into "confidence" and "critique" capacities, designing probabilistic metrics to evaluate them and exploring the role of reflection mechanisms in model behavior. Additionally, MCTSr [823], LLM2 Yang et al. [736], ReST-MCTS* [822] and ReARTeR [554] emphasize dynamic reflection through iterative error correction and confidence adjustments, allowing models to autonomously refine reasoning strategies [144]. He et al. [191] extend this paradigm to multi-agent scenarios, improving agent system performance [742]. Moreover, Yuksekgonul et al. [800] and Peng et al. [471] further expand the paradigm by enabling automatic prompt optimization driven by LLMs. This approach facilitates more generalized and automated refinement of input prompts across a range of tasks, as opposed to focusing solely on refining output results. However, without oracle feedback, RLLM’s self-refinement process fails, causing instability in both intermediate and final answers, leading to biases in simple factual queries and introducing cognitive biases in complex tasks [836, 721].
@@ 5.2.3 RL-based Refinement Learning
In recent research, several approaches have been proposed to enhance the performance of refinement through reinforcement learning. Earlier, Kumar et al. [286] observed that SFT of RLLMs often fails to promote self-refinement behaviors. This limitation stems from a distributional mismatch between data collection strategies and model responses, as well as the risk of behavioral collapse. To address this, SCoRe [286] enhances self-refinement by training the model on its own self-generated correction trajectories and employing regularization to guide the learning process. This method prioritizes fostering self-refinement during testing, rather than merely maximizing reward for specific prompts [810]. Further, Guo et al. [179] demonstrate that applying outcome-level rewarded RL can trigger an “Aha moment,” activating the model’s natural feedback and refinement behaviors without the need for human guidance. Moreover, Guo et al. [179], Zeng et al. [809] and Ma et al. [416] explore initializing LLMs with iterative self-verification and self-correction behaviors, which are strengthened through supervised fine-tuning and further enhanced by outcome-level RL. Ma et al. [416] and Yang et al. [741] extend these capabilities with process-level RL, minimizing resource usage while enabling adaptive reasoning refinements during inference. More recently, Lee et al. [300] introduce an intrinsic verifier module to decide when refinements should be applied, using RL to further encourage self-refinement when errors are detected.
@@ 6 Extensive Exploration for Long CoT
Current research focuses on several key areas: (1) Exploration Scaling (§ 6.1), which examines the breadth and depth of exploration and its effect on downstream applications; (2) Internal Exploration (§ 6.2), which emphasizes training models to develop internal exploration capabilities; and (3) External Exploration (§ 6.3), which investigates how models can leverage external systems to enhance their exploratory abilities.
Figure 9: Schematic representations of two common inference test-time scaling strategies: (a) vertical scaling, which extends the length of Long CoT but is constrained by the reasoning boundaries of RLLMs; and (b) parallel scaling, which increases the sample size and aggregates multiple outcomes, yet does not surpass the performance of Pass@k.
@@ 6.1 Exploration Scaling
Recent advances in inference-time scaling algorithms [262, 672, 50] have attracted significant interest, particularly in scaling reasoning length to improve performance [413, 448, 327]. Following Chen et al. [80], as shown in Figure 9, exploration scaling can be understood through two paradigms: (1) vertical scaling, akin to a series of resistors, which connects multiple reasoning processes using reflection; and parallel scaling, similar to parallel resistors, where a unified verification/feedback mechanism selects the most effective reasoning processes.
6.1.1 Vertical Scaling
Vertical scaling refers to extending the reasoning output within a single model generation, significantly boosting model performance [308, 837]. Early works by Fu et al. [146] and Jaech et al. [240] show that increasing the length of the reasoning path can greatly improve performance. Tian et al. [587] enhances model reasoning iteratively by using prior answers as prompts for each successive round, thus enabling vertical scaling of the reasoning process. Building on this, later studies [246, 315] further explore enhancing logical depth through tree-based searches within a fixed compute budget, resulting in notable performance gains. Building upon this, Muennighoff et al. [441] introduce a test-time scaling method that improves reasoning by fine-tuning and budget forcing, yielding substantial gains with additional computing at test time. To address the constraints of attention spans, some studies focus on expanding reasoning length in latent spaces. Geiping et al. [159] and Chen et al. [93] enhance test-time reasoning performance by implicitly scaling computation in latent space through recurrent depth.
@@ 6.1.2 Parallel Scaling
Parallel scaling refers to the process of increasing the number of reasoning iterations during model generation and then verfiy these results to get the final output, which significantly enhances model performance [2, 686, 50, 385]. Initially, Wang et al. [652] introduce the concept of self-consistency, demonstrating that multiple sampling processes followed by majority voting for effective exploration.
@@
The primary focus of recent research is optimizing verification, which can be categorized into two types: (1) Overall Verification: Recent works [887, 665] divide the scaling process into two stages: "reasoning" and "self-verification." By replacing majority voting in self-consistency with self-verification, these approaches show significant improvements [859, 70, 907, 291]. In code scenarios, WoT [851], CISC [566] and S* [316] scale the Long CoT in parallel, using output confidence or code execution results for verification, effectively assessing reasoning quality [505, 158, 217, 898]. Further, Nye et al. [449] and Weir et al. [671] train RLLMs to simulate code execution, removing the need for test cases in code-related parallel scaling. Chain-of-Verification [80] introduces meta-verification, sampling multiple verification instances to identify the correct one. Kim et al. [279], Chen et al. [95], and Vacareanu et al. [600] validate this approach empirically by evaluating answer correctness based on reasoning path properties. Moreover, Li et al. [340] tune a specific RLLM to verify and aggregate answers, showing improved performance. This suggests that PRM cannot replace a specially trained RLLM for verification due to training goal biases [856]. Finally, Kang et al. [270] leverage self-uncertainty to select the best results. (2) Step Verification: Building on this, numerous researchers have explored step-level or finer-grained verification [72, 366]. Notably, DIVERSE [343], SSC-CoT [868], and Fine-grained Self-Consistency [80] combine diverse reasoning paths with step-level verification. In addition, Snell et al. [536], Wu et al. [686], Luo et al. [407], Wang et al. [618], Wu et al. [679], and Liu et al. [386] investigate how optimal scaling strategies based on MCTS can enhance smaller language models’ performance. Their findings show that a 1B RLLM can outperform a 405B model on complex tasks through parallel scaling [783]. Despite these advancements in verification, Chen et al. [80] demonstrate that these strategies cannot surpass Best-of-N methods, suggesting that breakthroughs cannot solely rely on optimization-based verification [90].
@@
Another key area of research focuses on generating diverse but less paths or strategies for efficient scaling [691, 614, 69]. For instance, Zeng et al. [812] aggregate the shortest yet most varied reasoning paths for better scalability. Similarly, Du et al. [131] adjust the sampling temperature to increase diversity, leading to improved scaling. Zhang et al. [832] and Liu et al. [375] optimize both candidate solution generation (e.g., prompts, temperature, and top-p) and reward mechanisms (such as self-evaluation and reward types), offering diverse strategies for parallel scaling. Moreover, Qin et al. [490], Luo et al. [410], and Yu et al. [784] enhance RLLM reasoning by scaling sampling across multiple natural and programming languages or varied expressions. Finally, Yang et al. [747] introduces a method where a small set of seed data, with varied response lengths, guides the model to engage in deeper reasoning by selecting the shortest correct responses across various inference efforts.
Takeaways: Exploration Scaling Exploration Mechanisms in Long CoT Reasoning: Exploration strategies like hypothesis branching and error backtracking are vital for overcoming limitations in linear reasoning paths and enhancing model performance. Scaling Exploration: Exploration can be scaled through vertical and parallel strategies to improve reasoning depth and efficiency. Verification and Sampling Optimization: Refining verification techniques and optimizing sampling for diverse reasoning paths are key to improving exploration efficiency and performance in Long CoT tasks.

### v4

@@ 5.1.2 Process Feedback
Recent studies highlight the significance of feedback in developing effective PRMs for complex reasoning tasks, particularly in a step-level view [128, 402, 505]. (1) Process Annotated PRM Training: Earlier, Lightman et al. [426] demonstrate that training process feedback with human-annotated data (PRM800K) surpasses outcome supervision in creating reliable reward models. However, this approach requires significant human effort. To address this, Wang et al. [763] introduce Math-Shepherd, a dataset that generates step-by-step supervision using a Tree Search-inspired method [69, 963]. Following this, methods like QwQ [703], Skywork-o1 [544], AceMath [477], and PRIME [137] adopt similar techniques to enhance PRM performance. Additionally, Zhang et al. [998] propose entropy regularization to improve model convergence. Rather than focusing solely on the first error step, Full-Step-DPO [869] assigns rewards for the entire reasoning chain, including error steps. VersaPRM [977] extends PRMs across multiple domains, broadening their applicability. Similarly, Gu et al. [205] and Zhang et al. [1033] suggest training models with student preferences aligned to teacher preferences, ensuring effective preference distillation. Further, Wang et al. [777] propose VisualPRM400K and expand this paradigm to multimodal scenarios. (2) Outcome Annotated PRM Training: Alternative approaches, such as ReST-MCTS* [994], OVM [942], Implicit PRM [962], AutoPSV [483], and DVO [1000], leverage outcome supervision or implicit feedback to train PRMs, reducing the need for extensive human-annotated data [857, 617]. UAS [944] incorporates uncertainty-aware value models [258] into feedback predictions [472, 157, 910]. Additionally, Aurora [682] utilizes ensemble prompting strategies and reference answers for reverse verification, training stronger PRMs that better align with the Long CoT data distribution. Furthermore, PAV [624] suggests that rewards should reflect reasoning progress, as measured by changes in the likelihood of producing a correct future response before and after each step. Yang et al. [897], Lee et al. [357], Yoon et al. [938] extend these paradigms to the token level. Moreover, Chen et al. [105] expand these into interactive agent scenarios, allowing for automatically learning reward models from the environment without additional manual annotation. Wang et al. [802] equip a dual-layer MLP module to evaluate the reward at each step, successfully integrating the policy model and PRM into a unified interface without additional process annotations, reducing over 99% of PRM parameters for efficient reasoning.
@@
As PRM training remains heavily dependent on manually annotated data, recent research has explored methods for enabling models to generate their natural language feedback to optimize performance [876]. These approaches fall into two primary categories: (1) Model-Driven Feedback Reasoning: Earlier work such as React [920] and Reflexion [642] enhances RLLMs with natural language feedback at each action and reasoning step [183, 129, 84], improving decision-making in diverse tasks. Similarly, Step-DPO [346] uses RLLM to self-verify step-level positive and negative pairs for training through the DPO paradigm, achieving strong performance. Additionally, Sun et al. [674] propose a dynamic error classification framework that adapts based on model outputs, improving performance in mathematical reasoning tasks by addressing specific error patterns in math word problems. Furthermore, Xie et al. [855] and He et al. [231] iteratively apply MCTS to collect preference data, utilizing its forward-looking capabilities to decompose instance-level rewards into more precise step-level signals, thereby enhancing feedback accuracy. However, step-wise feedback often suffers from reliability issues, which can be mitigated by uncertainty quantification [936, 933], improving the reliability of step-wise verification in reward models for mathematical reasoning tasks. Moreover, Fu et al. [175] define the CoT Average Causal Effect (CACE) to capture causal relationships between steps, resulting in a causalized Long CoT where all steps are both correct and comprehensible. (2) Environment-Driven Feedback Reasoning: Given the increasing complexity of large models, there is growing interest in combining prompt-based LLMs with external environments to generate more interpretable and controllable feedback [851, 254]. For example, ORPS [958] and Drori et al. [152] minimize dependence on human annotations by using execution feedback, enabling models to autonomously refine their solutions. Additionally, Shrestha et al. [643] contribute by translating model outputs into Python code, helping to identify logical errors, gain insights into flawed reasoning processes, and guide improvements in mathematical reasoning. Xu et al. [863] integrate reasoning models with an interactive environment, enabling learning in more dynamic scenarios and creating a more generalizable self-learning framework.
@@ 5.2.1 Prompt-based Refinement Generation
Research on prompt-based refine generation focuses on enhancing the performance of LLMs through iterative self-refinement mechanisms [552, 1049, 93, 446, 990, 727, 788]. A prominent approach involves prompting RLLMs to generate initial outputs, followed by self-feedback that iteratively refines and improves performance across tasks such as dialogue generation and mathematical reasoning [619, 516, 1059, 642, 525, 326, 723, 459], which even much reduce the hallucinations [272, 296]. Noteworthy methods, like Self-Backtracking [909], Refiner [564], and BackMath [1015], allow LLMs to adjust their reasoning autonomously, reducing unnecessary complexity in decision-making [835]. Further, Havrilla et al. [224] extend the paradigm by integrating overall-level and step-level refinements, improving refinement performance. Yang et al. [915] propose a method to decompose the self-correction capability of LLMs into "confidence" and "critique" capacities, designing probabilistic metrics to evaluate them and exploring the role of reflection mechanisms in model behavior. Additionally, MCTSr [995], LLM2 [895], ReST-MCTS* [994] and ReARTeR [675] emphasize dynamic reflection through iterative error correction and confidence adjustments, allowing models to autonomously refine reasoning strategies [174]. He et al. [226] extend this paradigm to multi-agent scenarios, improving both reasoning and agent system performance [901, 1086]. Moreover, Yuksekgonul et al. [971] and Peng et al. [567] further expand the paradigm by enabling automatic prompt optimization driven by LLMs. This approach facilitates more generalized and automated refinement of input prompts across a range of tasks, as opposed to focusing solely on refining output results. However, without oracle feedback, RLLM’s self-refinement process fails, causing instability in both intermediate and final answers, leading to biases in simple factual queries and introducing cognitive biases in complex tasks [1011, 874].
@@ 5.2.3 RL-based Refinement Learning
In recent research, several approaches have been proposed to enhance the performance of refinement through reinforcement learning [646, 1016]. Earlier, Kumar et al. [339] observed that SFT of RLLMs often fails to promote self-refinement behaviors. This limitation stems from a distributional mismatch between data collection strategies and model responses, as well as the risk of behavioral collapse. To address this, SCoRe [339] enhances self-refinement by training the model on its own self-generated correction trajectories and employing regularization to guide the learning process. This method prioritizes fostering self-refinement during testing, rather than merely maximizing reward for specific prompts [980]. Further, Guo et al. [213] demonstrate that applying outcome-level rewarded RL can trigger an “Aha moment,” activating the model’s natural feedback and refinement behaviors without the need for human guidance. Moreover, Guo et al. [213], Zeng et al. [979] and Ma et al. [506] explore initializing LLMs with iterative self-verification and self-correction behaviors, which are strengthened through supervised fine-tuning and further enhanced by outcome-level RL. Ma et al. [506] and Yang et al. [900] extend these capabilities with process-level RL, minimizing resource usage while enabling adaptive reasoning refinements during inference. More recently, Lee et al. [355] introduce an intrinsic verifier module to decide when refinements should be applied, using RL to further encourage self-refinement when errors are detected.
@@ 6 Extensive Exploration for Long CoT
Current research focuses on three key areas: (1) Exploration Scaling (§ 6.1), which explores the breadth and depth of exploration and its impact on downstream applications, particularly in improving the size of the exploration path $m$ in Equation (3); (2) Internal Exploration (§ 6.2), which focuses on training models to develop internal exploration capabilities, enabling more efficient and effective generation of $m$ exploration paths $\{n_{i+j}\}^{m}_{j=1}$ in Equation (3); and (3) External Exploration (§ 6.3), which examines how models can leverage external systems to enhance their exploratory abilities, facilitating the selection of the most effective path $n_{i+j}$ from the $m$ exploration paths in Equation (3).
Figure 9: Schematic representations of two common inference-time scaling strategies: (a) sequential scaling, which extends the length of Long CoT but is constrained by the reasoning boundaries of RLLMs; and (b) parallel scaling, which increases the sample size and aggregates multiple outcomes, yet does not surpass the performance of Pass@k.
@@ 6.1 Exploration Scaling
Recent advances in inference-time scaling algorithms [314, 811, 53, 1013, 107] have attracted significant interest, particularly in scaling reasoning length to improve performance [501, 542, 385, 751]. Following Chen et al. [88], as shown in Figure 9, exploration scaling can be understood through two paradigms: (1) sequential scaling, akin to a series of resistors, which connects multiple reasoning processes using reflection; and parallel scaling, similar to parallel resistors, where a unified verification/feedback mechanism selects the most effective reasoning processes.
6.1.1 Sequential Scaling
Sequential scaling refers to extending the reasoning output within a single model generation, significantly boosting model performance [364, 1012, 329]. Early works by Fu et al. [177] and Jaech et al. [289] show that increasing the length of the reasoning path can greatly improve performance. Tian et al. [709] enhances model reasoning iteratively by using prior answers as prompts for each successive round, thus enabling sequential scaling of the reasoning process. Building on this, later studies [295, 371] further explore enhancing logical depth through tree-based searches within a fixed compute budget, resulting in notable performance gains [10, 588]. Building upon this, Muennighoff et al. [534] introduce a inference-time scaling method that improves reasoning by fine-tuning and budget forcing, yielding substantial gains with additional computing at inference time. To address the constraints of attention spans, some studies focus on expanding reasoning length in latent spaces. Geiping et al. [191] and Chen et al. [104] enhance inference-time reasoning performance by implicitly scaling computation in latent space through recurrent depth. Setlur et al. [626] identified three core aspects of sequential scaling: (1) linking skills to asymmetric capabilities in base LLMs, such as connecting easy verification with difficult exploration; (2) enhancing exploration in reinforcement learning by utilizing the “negative” gradient of error trajectories, which extends search paths and links additional asymmetries; and (3) creating dynamic exploration by aligning task difficulty with training token budgets through tailored curricula.
@@ 6.1.2 Parallel Scaling
Parallel scaling refers to the process of increasing the number of reasoning iterations during model generation and then verfiy these results to get the final output, which significantly enhances model performance [2, 831, 53, 462, 55, 1096]. Initially, Wang et al. [786] introduce the concept of self-consistency, demonstrating that multiple sampling processes followed by majority voting for effective exploration.
@@
The primary focus of recent research is optimizing verification, which can be categorized into two types: (1) Overall Verification: Recent works [1078, 801] divide the scaling process into two stages: "reasoning" and "self-verification." By replacing majority voting in self-consistency with self-verification, these approaches show significant improvements [1042, 77, 1106, 345, 405]. In code scenarios, WoT [1031], CISC [688] and S* [372] scale the Long CoT in parallel, using output confidence or code execution results for verification, effectively assessing reasoning quality [609, 190, 261, 1092]. Further, Nye et al. [543] and Weir et al. [810], Stoisser et al. [662] train RLLMs to simulate code execution, removing the need for test cases in code-related parallel scaling. Chain-of-Verification [88] introduces meta-verification, sampling multiple verification instances to identify the correct one. Kim et al. [332], Chen et al. [106], and Vacareanu et al. [723] validate this approach empirically by evaluating answer correctness based on reasoning path properties. Moreover, Li et al. [400] tune a specific RLLM to verify and aggregate answers, showing improved performance. This suggests that PRM cannot replace a specially trained RLLM for verification due to training goal biases [1037]. Finally, Kang et al. [322] leverage self-uncertainty to select the best results. (2) Step Verification: Building on this, numerous researchers have explored step-level or finer-grained verification [79, 437]. Notably, DIVERSE [404], SSC-CoT [1056], and Fine-grained Self-Consistency [88] combine diverse reasoning paths with step-level verification. In addition, a series of works [649, 831, 494, 742, 821, 463] try to investigate how optimal scaling strategies based on MCTS can enhance smaller language models’ performance. Their findings show that a 1B RLLM can outperform a 405B model on complex tasks through parallel scaling [951]. Despite these advancements in verification, Chen et al. [88] demonstrate that these strategies cannot surpass Best-of-N methods, suggesting that breakthroughs cannot solely rely on optimization-based verification [101].
@@
Another key area of research focuses on generating diverse but less paths or strategies for efficient scaling [838, 738, 76, 641, 421, 654]. For instance, Zeng et al. [982] aggregate the shortest yet most varied reasoning paths for better scalability. Similarly, Du et al. [154] adjust the sampling temperature to increase diversity, leading to improved scaling. Zhang et al. [1006] and Liu et al. [447] optimize both candidate solution generation (e.g., prompts, temperature, and top-p) and reward mechanisms (such as self-evaluation and reward types), offering diverse strategies for parallel scaling. Moreover, Qin et al. [591], Luo et al. [497], and Yu et al. [952] enhance RLLM reasoning by scaling sampling across multiple natural and programming languages or varied expressions. Finally, Yang et al. [908] introduces a method where a small set of seed data, with varied response lengths, guides the model to engage in deeper reasoning by selecting the shortest correct responses across various inference efforts.
Takeaways: Exploration Scaling Exploration Mechanisms in Long CoT Reasoning: Exploration strategies like hypothesis branching and error backtracking are vital for overcoming limitations in linear reasoning paths and enhancing model performance. Scaling Exploration: Exploration can be scaled through sequential and parallel strategies to improve reasoning depth and efficiency. Verification and Sampling Optimization: Refining verification techniques and optimizing sampling for diverse reasoning paths are key to improving exploration efficiency and performance in Long CoT tasks.